Ergonomics Inc. sells ergonomically designed office chairs. The company has the following information: Average demand = 28 units per day Average lead time = 43 days Item unit cost = $63 for orders of less than 330 units Item unit cost = $62 for orders of 330 units or more Ordering cost = $38 Inventory carrying cost = 20% The business year is 250 days Assume there is no uncertainty at all about the demand or the lead time.

a. Calculate EOQ if unit cost is $63 and $62. (Note: These EOQs do not need to be feasible in their price range.) (Round up your answers to the next whole number.)

b. Calculate annual ordering costs for each alternative? (Round your answers to 2 decimal places.)

c. Calculate annual inventory carrying costs for each alternative? (Round your answers to 2 decimal places.)

d. Calculate annual product costs for each alternative?
Answer #1

Answers are in BOLD

demand per day 28 28
Days per year= 250 250
Annual demand, D= demand per day*days per year 7000 7000
cost per order, S $38.00 $38.00
item cost $63.00 $62.00
holding cost percent 20% 20.0%
holding cost per year, H= holding cost percent*item cost $12.60 $12.40
Lead time, LT 43 43
Answer a Q= SQRT(2*S*D/H) 205 207
#orders per year = D/Q 34.07 33.80
Answer c annual holding cost= Q*H/2 $1,294.53 $1,284.21
Answer b Annual ordering cost= No. of orders*order cost $1,294.53 $1,284.21
Answer d Purchasing cost= annual demand*item cost $441,000 $434,000

Please ask, if you have any doubts through the comment section. Do rate the answer